OpenStudy (anonymous):

If s denotes the length of the arc of a circle of radius r subtended by a central angle \theta , find the missing quantity. s = 7.02 meters, \theta = 2.6 radians, r = ? a. 0.37 m b. 3.6 m c. 1.35 m d. 2.7 m

OpenStudy (anonymous):

It will be 0.37 meters. The answer comes from the definition of radians: \[v= \frac{s}{r}\] When solving for r gives you 0.37m with the given radian (v), the given arc length (s).

OpenStudy (anonymous):

I think it will be 2.7 m What do you think?

OpenStudy (anonymous):

Are you sure about the your answer Elek?

OpenStudy (anonymous):

No ofc what the hell am I doing. It is wrong. Haha, it is r = s/v, I am so sorry

OpenStudy (anonymous):

Which gives 2.7. Again, deeply sorry. It's been a messy day^^

OpenStudy (anonymous):

Question 15 Perform the indicated operations and simplify the result. Leave the answer in factored form. \LARGE \frac{\frac{4}{x} + \frac{7}{x^2}}{\frac{16}{x^2} - \frac{49}{x}} a. \frac{1}{4 -7x} b. \frac{4x^2 + 7}{16 - 49x} c. \frac{4x + 7}{16 - 49x} d. \frac{1}{4x - 7}

OpenStudy (anonymous):

girl : Am I pretty? Boy : NO. Girl : Do you want to be with me forever? Boy : NO. Girl : Would you cry if I walked away?Boy : NO. She heard enough, and was hurt. She walked away, tears ran down her face. The boy grabbed her arm. Boy : Your not pretty, your beautiful. I don't want to be with you forever, I NEED to be with you forever. And I wouldn't cry if you walked away, I would DIE. (Boy whispers) : Please? Stay with me. (Girl whispers) : I will. Tonight at midnight your true love will realize they love you. Something good will happen to you between 1-4 pm. Tomorrow it could be anywhere. Get ready for the biggest shock of your life! If you don't post this to 5 other pages. You will have relationship problems for the next 10 years